# Density-Aware Farthest Point Sampling

### v1

@@ Appendix D Asymptotic unbiasedness of the density estimations
To show that $\lim\limits_{b\rightarrow\infty}r_{b,n}^{k}(\bar{{\bm{x}}})=0$ it is sufficient to observe that $\lim\limits_{b\rightarrow\infty}\min\limits_{\bar{{\bm{x}}}_{j}\in\mathcal{L}_{b}}\|\tilde{{\bm{x}}}-\bar{{\bm{x}}}_{j}\|_{2}=0$ due to the fact that as the number of samples in $\mathcal{L}_{b}$ increases, the distance between $\bar{{\bm{x}}}$ and its nearest neighbor in $\mathcal{L}_{b}$ decreases, tending to zero. Moreover, $\lim\limits_{b\rightarrow\infty}\frac{\epsilon_{\mathcal{X}}}{b}=0$. Similarly, we have that $\lim\limits_{n\rightarrow\infty}r_{b,n}^{k}({\bm{x}})=0$ from the fact that $\lim\limits_{n\rightarrow\infty}\rho_{k,n}({\bm{x}})=0$ which follows from the observation that as the number of samples in $\mathcal{D}_{n}$ increases, the distance between $\bm{x}$} and its $k$-nearest neighbor decreases, tending to zero. Thus, we can apply Theorem D.1 to $\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}_{n}}}({\bm{x}})$ and $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}_{b}}}(\bar{{\bm{x}}})$ showing that they are asymptotically unbiased estimators of $p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$ and $p_{\mathcal{X}_{\mathcal{L}}}(\bar{{\bm{x}}})$, respectively, for any $2\leq k<n$.
@@ Appendix F Computational efficiency DA-FPS
Given the novelty of DA-FPS, it is important to investigate its computational complexity. DA-FPS can be implemented using $\mathcal{O}(|\mathcal{D}|k)$ memory and the greedy selection takes $\mathcal{O}(db|\mathcal{D}|k)$ time. $|\mathcal{D}|$ is the amount of available data points, $k$ the amount of nearest neighbors we consider for the density approximation, $b$ is the amount of points we select and $d$ is the dimension of the data points. The computational cost of DA-FPS is determined by the weights update (line 7 in Algorithm 1) taking $\mathcal{O}(d|\mathcal{D}|k)$ at each of the $b$ iterations. In the current implementation the weights update involves iterating over all points in $\mathcal{D}$ and compute the distances between the new selected point and the points’ $k$-nearest neighbors, which costs $\mathcal{O}(d|\mathcal{D}|k)$. Note that, initializing DA-FPS requires the computation of the $k$-nearest neighbors matrix, which can be a potential bottleneck. In our implementation we query the $k$-nearest neighbors using the cKDtree algorithm from the SciPy python library (Virtanen et al., 2020). The algorithm takes $\mathcal{O}(d|\mathcal{D}|\log|\mathcal{D}|)$ for building the balanced tree in the worst case scenario. After that, it queries the $k$-nearest neighbors with a worst-case cost of $\mathcal{O}(|\mathcal{D}|^{1-\frac{1}{d}})$ and an average cost of $\mathcal{O}(\log|\mathcal{D}|)$. Additionally, it is important to note that if the nearest neighborhood size to compute the weights as in (8) is set to $k=1$, the optimization problem in (10) coincides with the fill distance minimization problem and Algorithm 1 reduces to the well known Farthest Point Sampling algorithm (FPS), thus providing $2$-optimal solution (Har-Peled, 2011)., which is the best approximation factor attainable in polynomial time with theoretical guarantees (Hochbaum & Shmoys, 1985).
We implemented two versions of DA-FPS: one using NumPy (van der Walt et al., 2011) and the other with PyTorch (Paszke et al., 2019). The average computation times (over five runs) to select 20% of data points from QM7, QM8, QM9, and ZINC are respectively as follows: NumPy - 6, 48, 1974, and 70 seconds; PyTorch - 4, 31, 968, and 33 seconds. This was conducted on a 48-core CPU with 384 GB RAM. DA-FPS was initialized with $u=0$ and $k=100$ independently of the dataset. DA-FPS’ PyTorch implementation is faster than the NumPy implementation. This is because PyTorch can run computations exploiting multiple CPU cores. It uses libraries like OpenMP(Dagum & Menon, 1998) and Math Kernel Library (Wang et al., 2014) to perform operations on multiple CPU cores, leading to faster computations.
@@ Appendix G Datasets
This appendix draws from (Climaco & Garcke, 2024) and provides additional information related to the datasets, preprocessing procedures and molecular descriptors used for the experiments reported in Section 8
QM7 (Blum & Reymond, 2009; Rupp et al., 2012) contains 7,165 small organic molecules (up to 23 atoms, including C, N, O, S). Each molecule is represented by the upper triangular entries of its Coulomb matrix (Rupp et al., 2012), resulting in a feature vector in $\mathbb{R}^{276}$. The regression target is the atomization energy (in eV), i.e., the energy required to separate all atoms in a molecule.
@@
QM9 (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) consists of 130,202 small organic molecules with up to 9 heavy atoms. For each molecule it provides its SMILES string and 19 computed properties. Following along (Climaco & Garcke, 2024) we preprocess QM9 by removing molecules that fail consistency checks, cannot be parsed by RDKit (Landrum, 2012), or have duplicate SMILES. Molecular features are computed using Mordred (Moriwaki et al., 2018), with missing values set to zero, features with zero variance removed, normalization to (0, 1), and PCA to 100 dimensions. The regression target is the HOMO-LUMO energy gap (in eV), an important indicator of molecular stability.
The ZINC dataset (Gómez-Bombarelli et al., 2018) consists of about 250,000 molecules with up to 38 heavy atoms selected from the ZINC database (Sterling & Irwin, 2015), which contains over 120 million purchasable organic molecules. To reduce the computational effort of our analysis, we follow along Dwivedi et al. (2023) and consider a subset of the ZINC dataset. Specifically, we use a subset of ZINC consisting of 24000 molecules selected uniformly at random. The molecular representation we employ is based on the Mordred (Moriwaki et al., 2018) library, as for the QM8 and QM9 datasets. We normalize the features provided by the Mordred library, to scale them independently in the interval (0, 1). Again, we set to zero all the descriptor values that Mordred could not compute, removed the features for which the values across the dataset have zero variance and applied PCA to reduce the dimension of the feature vectors to 100. The label value to predict is the water-octanal partition coefficient (LogP), describing the molecules’ solubility.
@@ G.1 Datasets for additional experiments in Appendix K
This section provides a more detailed description of the additional datasets unrelated to quantum chemistry used for experiments in Appendix K, including information on the preprocessing procedures.
The Concrete Compressive Strength dataset (Yeh, 1998) downloaded from the UCI Machine Learning Repository (Dua & Graff, 2017) contains 1030 data points and is used for regression tasks. It includes eight features: the amounts of cement, blast furnace slag, fly ash, water, superplasticizer, coarse aggregate, and fine aggregate, as well as the age of the concrete in days. We remove 34 data points having identical descriptors as at least one other point in the dataset, obtaining a reduced dataset of 996 data points. Furthermore, we normalize the features to scale them independently in the interval (0, 1). The target variable is the compressive strength of the concrete, measured in megapascals (MPa). This dataset is used to test machine learning models to predict material properties.
In the additional experiments illustrated in Appendix K, we use the Twinning algorithm implementation from Vakayil & Joseph (2022), which selects subsets based on an integer $r$, the inverse of the partitioning ratio. Since the algorithm strictly partitions the dataset according to this ratio, we remove 6 points from the Concrete dataset, resulting in a reduced dataset with 990 points. The points were selected randomly. This adjustment ensures that the subset size determined by the Twinning algorithm matches the percentages used to select the subsets, eliminating any discrepancies. Similarly, for the experiments in Appendix K, we remove 66 points from the QM8 dataset, creating a reduced dataset of 21700 points. QM8 is preprocessed with the same procedure used in Section 8.
The Electrical Grid Stability Simulated dataset from the UCI Machine Learning Repository (Dua & Graff, 2017) contains 10000 data points and is designed for both classification and regression tasks. Each data point in this dataset is represented by 12 features that describe characteristics of a simulated power grid. We normalize the features to scale them independently in the interval (0, 1). For regression tasks, the target variable is the stability margin, which quantifies the power grid’s stability.
@@ Appendix H Data assumptions
For the QM datasets, the data assumptions have been already tested in Climaco & Garcke (2024) where the authors perform experiments related to FPS. However, for the experiments with DA-FPS we use different feature vectors. Thus, it is worth to verify whether the new feature vectors we consider still respect the required assumptions or not. We use the same procedure employed in Climaco & Garcke (2024) and study the correlation between pairwise distances in the feature and labels spaces by computing the Pearson’s $(\rho_{p})$ and Spearman’s $(\rho_{s})$ correlation coefficients. We recall that these coefficients measure and quantify the correlation between the quantities of interest. We compute the correlation coefficients for the pairwise distances in the feature and label spaces on the QM7, QM8, QM9 and ZINC. We consider the features and labels used for the experiments illustrated in Section 8 and Appendix M. Due to memory issue in storing the distance matrix, for the QM9 we computed the correlation coefficients on 50% of randomly selected data points. We selected random subsets and computed the associated correlation coefficients for five times. The results we report for the QM9 are the average over the five runs. The computed coefficients are 0.15, 0.22, 0.26 and 0.22 for $\rho_{p}$, and 0.27, 0.19, 0.22 and 0.19 $\rho_{s}, for$ QM7, QM8, QM9 and ZINC, respectively. Thus, both the Pearson’s and Spearman’s coefficients indicate a positive correlation between the pairwise distances of the data features and labels, suggesting that the data assumption considered in Theorem 4.2 is respected for each of the considered datasets.
@@ Appendix I Regression models: KRR and FNN
In this work, we follow along Climaco & Garcke (2024) and use two regression models already used in previous works for molecular property prediction: kernel ridge regression with a Gaussian kernel (KRR) and feed-forward neural networks (FNN). In this Appendix we recapture the key aspects of the KRR and FNN regression models. For a more in dept analysis of KRR and FNN for molecular property prediction see Deringer et al. (2021) and Pinheiro et al. (2020), respectively.
@@
In this appendix we use the root mean squared error (RMSE) to evaluate the performance of DA-FPS against the baselines used in Section 8. Given true target values $\{y_{i}\}_{i=1}^{n_{u}}$ and the predicted values $\{\tilde{y}_{i}\}_{i=1}^{n_{u}}$ the RMSE is defined as $\text{RMSE}:=\sqrt{\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}|y_{i}-\tilde{y}_{i}|^{2}}$, where $n_{u}$ is the number of unlabeled points in the data pool. Similarly to the MAE, the RMSE provides information on the average quality of the prediction. The difference is in the fact that RMSE penalizes large errors more than the MAE, thus providing insight on the robustness of the predictions. The graphs in Fig. 3 compare the performances of DA-FPS with the baselines and illustrate the RMSE of the predictions for the KRR (top row) and FNN (bottom row). The illustrated results suggest that the performances of FPS may be closer to that of DA-FPS when we consider the RMSE. This is particularly evident on the smaller QM7 and for larger set sizes on the QM8. Note that the RMSE penalizes large errors more than the MAE, thus giving more relevance to outlier error values. As we know from Climaco & Garcke (2024) FPS leads to a substantial decrease in maximum prediction error and hence reduces the amount and magnitude of large error values. Nevertheless, DA-FPS still leads competitive performances in terms of the RMSE, highlighting its robustness against large errors.
@@
In this section we present experiments considering two additional datasets unrelated to quantum chemistry, two additional sampling strategies, and one additional kernel method. The new datasets are the Concrete Compressive Strength dataset and the Electrical Grid Stability Simulated Dataset, from the public UCI ML repository (Dua & Graff, 2017). The Electrical Grid Stability dataset contains 10000 data points represented by 12-dimensional vectors. The target variable for regression is the stability margin (stab), reflecting grid stability. The Concrete dataset consists of 1030 data points of described by 8-dimensional vectors. The target variable is the compressive strength in megapascals (MPa). In Section G.1 we provide additional details on the datasets and preprocessing procedures. We also consider the QM8 to evaluate the performance of the additional sampling strategies and regression model in a quantum chemistry context.
The additional sampling strategies are the Twinning algorithm (Vakayil & Joseph, 2022), and the facility location with a Gaussian similarity function (FacLocG) as defined in Bhatt et al. (2024). Fine-tuning is required for the Gaussian width of the similarity function in facility location, and we follow the methodology outlined in Bhatt et al. (2024), with complete details and hyperparameters described in Section M.1 of this appendix. Section M.1 also includes the hyperparameters used for DA-FPS. In these experiments, we compare DA-FPS with the additional sampling strategies alongside the baseline methods used in the previous section, including RDM, $k$-medoids++, facility location and FPS. The Twinning algorithm implementation from Vakayil & Joseph (2022) only allows the selection of subsets of the size that can be expressed as an integer “rr” representing the inverse of the partitioning ratio, i.e., for obtaining a training subset consisting of 20% of the data points, we must set $r=\frac{100}{20}=5$. Consequently, to use the Twinning algorithm, we consider training set sizes similar to those considered in the previous section, but that can be expressed as an integer ratio. Specifically, we consider training set sizes of 5%, 10%, 16.67% and 20% associated with a ratio $r$ of 20, 10, 6 and 5, respectively.
@@
Theorem 7.1 states that DA-FPS provides a $2k$-optimal result to the optimization problem in (10), which considers dynamic-weights, that is, the weights are iteratively updated any time a new point is selected. We can consider a simplified version of the optimization problem in (10) by considering for each data point ${\bm{x}}\in\mathcal{D}_{\mathcal{X}}$ a fixed weight $\omega({\bm{x}})\in\mathbb{R}^{+}$, which is determined a-priori and does not depend on the selected set. Such a simplified scenario has been already studied in literature. In particular, the authors of Dyer & Frieze (1985) show that it is possible to find solutions that are $\sigma$-optimal, with $\sigma:=\min\{3,1+\alpha\}$, where $\alpha:=\frac{\max_{{\bm{x}}\in\mathcal{D}_{\mathcal{X}}}\omega({\bm{x}})}{\min_{{\bm{x}}\in\mathcal{D}_{\mathcal{X}}}\omega({\bm{x}})}$. That is, it is possible to find solutions that are at least 3-optimal. With the following theorem we attempt to extend the result provided in Dyer & Frieze (1985) into our scenario with dynamic weights.
@@
$\sigma:=\min\{3,1+\alpha\}\enspace\text{ with }\enspace\alpha:=\max_{\begin{subarray}{c}i,j=1,\dots,b+1\\ i<j\end{subarray}}\frac{\omega^{k}_{\mathcal{L}_{j-1}}({\bm{x}}_{j})}{\omega^{k}_{\mathcal{L}_{i-1}}({\bm{x}}_{i})}.$ (36)

### v2

@@ Appendix D Asymptotic unbiasedness of the density estimations
To show that $\lim\limits_{b\rightarrow\infty}r_{b,n}^{k}(\bar{{\bm{x}}})=0$ it is sufficient to observe that $\lim\limits_{b\rightarrow\infty}\min\limits_{\bar{{\bm{x}}}_{j}\in\mathcal{L}_{b}}\|\tilde{{\bm{x}}}-\bar{{\bm{x}}}_{j}\|_{2}=0$ because as the number of samples in $\mathcal{L}_{b}$ increases, the distance between $\bar{{\bm{x}}}$ and its nearest neighbor in $\mathcal{L}_{b}$ decreases, tending to zero. Moreover, $\lim\limits_{b\rightarrow\infty}\frac{\epsilon_{\mathcal{X}}}{b}=0$. Similarly, we have that $\lim\limits_{n\rightarrow\infty}r_{b,n}^{k}({\bm{x}})=0$ from the fact that $\lim\limits_{n\rightarrow\infty}\rho_{k,n}({\bm{x}})=0$ which follows from the observation that as the number of samples in $\mathcal{D}_{n}$ increases, the distance between $\bm{x}$} and its $k$-nearest neighbor decreases, tending to zero. Thus, we can apply Theorem D.1 to $\hat{p}^{k}_{\mathcal{X}_{\mathcal{D}_{n}}}({\bm{x}})$ and $\hat{p}^{k}_{\mathcal{X}_{\mathcal{L}_{b}}}(\bar{{\bm{x}}})$ showing that they are asymptotically unbiased estimators of $p_{\mathcal{X}_{\mathcal{D}}}({\bm{x}})$ and $p_{\mathcal{X}_{\mathcal{L}}}(\bar{{\bm{x}}})$, respectively, for any $2\leq k<n$.
@@ Appendix F Computational efficiency DA-FPS
Given the novelty of DA-FPS, it is important to investigate its computational complexity. DA-FPS can be implemented using $\mathcal{O}(|\mathcal{D}|k)$ memory and the greedy selection takes $\mathcal{O}(db|\mathcal{D}|k)$ time. $|\mathcal{D}|$ is the amount of available data points, $k$ the amount of nearest neighbors we consider for the density approximation, $b$ is the amount of points we select and $d$ is the dimension of the data points. The computational cost of DA-FPS is determined by the weights update (line 7 in Algorithm 1) taking $\mathcal{O}(d|\mathcal{D}|k)$ at each of the $b$ iterations. In the current implementation the weights update involves iterating over all points in $\mathcal{D}$ and compute the distances between the new selected point and the points’ $k$-nearest neighbors, which costs $\mathcal{O}(d|\mathcal{D}|k)$. Note that, initializing DA-FPS requires the computation of the $k$-nearest neighbors matrix, which can be a potential bottleneck. In our implementation we query the $k$-nearest neighbors using the cKDtree algorithm from the SciPy python library (Virtanen et al., 2020). The algorithm takes $\mathcal{O}(d|\mathcal{D}|\log|\mathcal{D}|)$ for building the balanced tree in the worst case scenario. After that, it queries the $k$-nearest neighbors with a worst-case cost of $\mathcal{O}(|\mathcal{D}|^{1-\frac{1}{d}})$ and an average cost of $\mathcal{O}(\log|\mathcal{D}|)$. Additionally, it is important to note that if the nearest neighborhood size to compute the weights as in (8) is set to $k=1$, the optimization problem in (10) coincides with the fill distance minimization problem and Algorithm 1 reduces to the well known Farthest Point Sampling algorithm (FPS), thus providing $2$-optimal solution (Har-Peled, 2011)., which is the best approximation factor attainable in polynomial time with theoretical guarantees (Hochbaum and Shmoys, 1985).
We implemented two versions of DA-FPS: one using NumPy (van der Walt et al., 2011) and the other with PyTorch (Paszke et al., 2019). The average computation times (over five runs) to select 20% of data points from QM7, QM8, QM9, and ZINC are respectively as follows: NumPy - 6, 48, 1974, and 70 seconds; PyTorch - 4, 31, 968, and 33 seconds. This was conducted on a 48-core CPU with 384 GB RAM. DA-FPS was initialized with $u=0$ and $k=100$ independently of the dataset. DA-FPS’ PyTorch implementation is faster than the NumPy implementation. This is because PyTorch can run computations exploiting multiple CPU cores. It uses libraries like OpenMP(Dagum and Menon, 1998) and Math Kernel Library (Wang et al., 2014) to perform operations on multiple CPU cores, leading to faster computations.
@@ Appendix G Datasets
This appendix draws from (Climaco and Garcke, 2024) and provides additional information related to the datasets, preprocessing procedures and molecular descriptors used for the experiments reported in Section 8
QM7 (Blum and Reymond, 2009; Rupp et al., 2012) contains 7,165 small organic molecules (up to 23 atoms, including C, N, O, S). Each molecule is represented by the upper triangular entries of its Coulomb matrix (Rupp et al., 2012), resulting in a feature vector in $\mathbb{R}^{276}$. The regression target is the atomization energy (in eV), i.e., the energy required to separate all atoms in a molecule.
@@
QM9 (Ruddigkeit et al., 2012; Ramakrishnan et al., 2014) consists of 130,202 small organic molecules with up to 9 heavy atoms. For each molecule it provides its SMILES string and 19 computed properties. Following along (Climaco and Garcke, 2024) we preprocess QM9 by removing molecules that fail consistency checks, cannot be parsed by RDKit (Landrum, 2012), or have duplicate SMILES. Molecular features are computed using Mordred (Moriwaki et al., 2018), with missing values set to zero, features with zero variance removed, normalization to (0, 1), and PCA to 100 dimensions. The regression target is the HOMO-LUMO energy gap (in eV), an important indicator of molecular stability.
The ZINC dataset (Gómez-Bombarelli et al., 2018) consists of about 250,000 molecules with up to 38 heavy atoms selected from the ZINC database (Sterling and Irwin, 2015), which contains over 120 million purchasable organic molecules. To reduce the computational effort of our analysis, we follow along Dwivedi et al. (2023) and consider a subset of the ZINC dataset. Specifically, we use a subset of ZINC consisting of 24000 molecules selected uniformly at random. The molecular representation we employ is based on the Mordred (Moriwaki et al., 2018) library, as for the QM8 and QM9 datasets. We normalize the features provided by the Mordred library, to scale them independently in the interval (0, 1). Again, we set to zero all the descriptor values that Mordred could not compute, removed the features for which the values across the dataset have zero variance and applied PCA to reduce the dimension of the feature vectors to 100. The label value to predict is the water-octanal partition coefficient (LogP), describing the molecules’ solubility.
@@ G.1 Datasets for additional experiments in Appendix K
In this section we provide a more detailed description of the additional datasets unrelated to quantum chemistry used for experiments in Appendix K, including information on the preprocessing procedures.
The Concrete Compressive Strength dataset (Yeh, 1998) downloaded from the UCI Machine Learning Repository (Dua and Graff, 2017) contains 1030 data points and is used for regression tasks. It includes eight features: the amount of cement, blast furnace slag, fly ash, water, superplasticizer, coarse aggregate, and fine aggregate, as well as the age of the concrete in days. We remove 34 data points having identical descriptors as at least one other point in the dataset, obtaining a reduced dataset of 996 data points. Furthermore, we normalize the features to scale them independently in the interval $(0,1)$. The target variable is the compressive strength of the concrete, measured in megapascals (MPa). This dataset is used to test machine learning models to predict material properties.
In the additional experiments illustrated in Appendix K, we use the Twinning algorithm implementation from Vakayil and Joseph (2022), which selects subsets based on an integer $r$, the inverse of the partitioning ratio. Since the algorithm strictly partitions the dataset according to this ratio, we remove 6 points from the Concrete dataset, resulting in a reduced dataset with 990 points. The points were selected randomly. This adjustment ensures that the subset size determined by the Twinning algorithm matches the percentages used to select the subsets, eliminating any discrepancies. Similarly, for the experiments in Appendix K, we remove 66 points from the QM8 dataset, creating a reduced dataset of 21700 points. QM8 is preprocessed with the same procedure used in Section 8.
The Electrical Grid Stability Simulated dataset from the UCI Machine Learning Repository (Dua and Graff, 2017) contains 10000 data points and is designed for both classification and regression tasks. Each data point in this dataset is represented by 12 features that describe characteristics of a simulated power grid. We normalize the features to scale them independently in the interval $(0,1)$. For regression tasks, the target variable is the stability margin, which quantifies the power grid’s stability.
@@ Appendix H Data assumptions
For the QM datasets, the data assumptions have been already tested in Climaco and Garcke (2024) where the authors perform experiments related to FPS. However, for the experiments with DA-FPS we use different feature vectors. Thus, it is worth to verify whether the new feature vectors we consider still respect the required assumptions or not. We use the same procedure employed in Climaco and Garcke (2024) and study the correlation between pairwise distances in the feature and labels spaces by computing the Pearson’s $(\rho_{p})$ and Spearman’s $(\rho_{s})$ correlation coefficients. We recall that these coefficients measure and quantify the correlation between the quantities of interest. We compute the correlation coefficients for the pairwise distances in the feature and label spaces on the QM7, QM8, QM9 and ZINC. We consider the features and labels used for the experiments illustrated in Section 8 and Appendix M. Due to memory issue in storing the distance matrix, for the QM9 we computed the correlation coefficients on 50% of randomly selected data points. We selected random subsets and computed the associated correlation coefficients for five times. The results we report for the QM9 are the average over the five runs. The computed coefficients are 0.15, 0.22, 0.26 and 0.22 for $\rho_{p}$, and 0.27, 0.19, 0.22 and 0.19 $\rho_{s}, for$ QM7, QM8, QM9 and ZINC, respectively. Thus, both the Pearson’s and Spearman’s coefficients indicate a positive correlation between the pairwise distances of the data features and labels, suggesting that the data assumption considered in Theorem 4.2 is respected for each of the considered datasets.
@@ Appendix I Regression models: KRR and FNN
In this work, we follow along Climaco and Garcke (2024) and use two regression models already used in previous works for molecular property prediction: kernel ridge regression with a Gaussian kernel (KRR) and feed-forward neural networks (FNN). In this appendix we recapture the key aspects of the KRR and FNN regression models. For a more in depth analysis of KRR and FNN for molecular property prediction see Deringer et al. (2021) and Pinheiro et al. (2020), respectively.
@@
In this appendix we use the root mean squared error (RMSE) to evaluate the performance of DA-FPS against the baselines used in Section 8. Given true target values $\{y_{i}\}_{i=1}^{n_{u}}$ and the predicted values $\{\tilde{y}_{i}\}_{i=1}^{n_{u}}$ the RMSE is defined as $\text{RMSE}:=\sqrt{\frac{1}{n_{u}}\sum_{i=1}^{n_{u}}|y_{i}-\tilde{y}_{i}|^{2}}$, where $n_{u}$ is the number of unlabeled points in the data pool. Similarly to the MAE, the RMSE provides information on the average quality of the prediction. The difference is in the fact that RMSE penalizes large errors more than the MAE, thus providing insight on the robustness of the predictions. The graphs in Fig. 3 compare the performances of DA-FPS with the baselines and illustrate the RMSE of the predictions for the KRR (top row) and FNN (bottom row). The illustrated results suggest that the performances of FPS may be closer to that of DA-FPS when we consider the RMSE. This is particularly evident on the smaller QM7 and for larger set sizes on the QM8. Note that the RMSE penalizes large errors more than the MAE, thus giving more relevance to outlier error values. As we know from Climaco and Garcke (2024), FPS leads to a substantial decrease in maximum prediction error and hence reduces the amount and magnitude of large error values. Nevertheless, DA-FPS still leads to competitive performances in terms of the RMSE, highlighting its robustness against large errors.
@@
In this section we present experiments considering two additional datasets unrelated to quantum chemistry, two additional sampling strategies, and one additional kernel method. The new datasets are the Concrete Compressive Strength dataset and the Electrical Grid Stability Simulated Dataset, from the public UCI ML repository (Dua and Graff, 2017). The Electrical Grid Stability dataset contains 10000 data points represented by 12-dimensional vectors. The target variable for regression is the stability margin (stab), reflecting grid stability. The Concrete dataset consists of 1030 data points described by 8-dimensional vectors. The target variable is the compressive strength in megapascals (MPa). In Section G.1 we provide additional details on the datasets and preprocessing procedures. We also consider the QM8 to evaluate the performance of the additional sampling strategies and regression model in a quantum chemistry context.
The additional sampling strategies are the Twinning algorithm (Vakayil and Joseph, 2022), and the facility location with a Gaussian similarity function (FacLocG) as defined in Bhatt et al. (2024). Fine-tuning is required for the Gaussian width of the similarity function in facility location, and we follow the methodology outlined in Bhatt et al. (2024), with complete details and hyperparameters described in Section M.1 of this appendix. Section M.1 also includes the hyperparameters used for DA-FPS. In these experiments, we compare DA-FPS with the additional sampling strategies alongside the baseline methods used in the previous section, including RDM, $k$-medoids++, facility location and FPS. The Twinning algorithm implementation from Vakayil and Joseph (2022) only allows the selection of subsets of the size that can be expressed as an integer “rr” representing the inverse of the partitioning ratio, i.e., for obtaining a training subset consisting of 20% of the data points, we must set $r=\frac{100}{20}=5$. Consequently, to use the Twinning algorithm, we consider training set sizes similar to those considered in the previous section, but that can be expressed as an integer ratio. Specifically, we consider training set sizes of 5%, 10%, 16.67% and 20% associated with a ratio $r$ of 20, 10, 6 and 5, respectively.
@@
Theorem 7.1 states that DA-FPS provides a $2k$-optimal result to the optimization problem in (10), which considers dynamic-weights, that is, the weights are iteratively updated any time a new point is selected. We can consider a simplified version of the optimization problem in (10) by considering for each data point ${\bm{x}}\in\mathcal{D}_{\mathcal{X}}$ a fixed weight $\omega({\bm{x}})\in\mathbb{R}^{+}$, which is determined a-priori and does not depend on the selected set. Such a simplified scenario has been already studied in literature. In particular, the authors of Dyer and Frieze (1985) show that it is possible to find solutions that are $\sigma$-optimal, with $\sigma:=\min\{3,1+\alpha\}$, where $\alpha:=\frac{\max_{{\bm{x}}\in\mathcal{D}_{\mathcal{X}}}\omega({\bm{x}})}{\min_{{\bm{x}}\in\mathcal{D}_{\mathcal{X}}}\omega({\bm{x}})}$. That is, it is possible to find solutions that are at least 3-optimal. With the following theorem we attempt to extend the result provided in Dyer and Frieze (1985) into our scenario with dynamic weights.
@@
$\sigma:=\min\{3,1+\alpha\}\kern 5.0pt\text{ with }\kern 5.0pt\alpha:=\max_{\begin{subarray}{c}i,j=1,\dots,b+1\\ i<j\end{subarray}}\frac{\omega^{k}_{\mathcal{L}_{j-1}}({\bm{x}}_{j})}{\omega^{k}_{\mathcal{L}_{i-1}}({\bm{x}}_{i})}.$ (36)